OpenStudy (anonymous):

How do you solve: 3x/x+5=-2? ( /=divided,over)

OpenStudy (anonymous):

3x/(x+5) = -2 3x = -2(x+5) 3x = -2x - 10 3x+2x = -10 5x = -10 x = -10/5 x = -2

OpenStudy (anonymous):

Ty.